UNIT COSTS, MULTIPLE PRODUCTS, VARIANCE ANALYSIS, SERVICE SETTING

The maternity wing of the city hospital has two types of patients: normal and caesarean. The standard quantities of labour and materials per delivery for 2011 are:

Normal           Caesarean

Direct materials (kg)                             8                       20

Nursing labour (hrs.)                           2                         4

The standard price paid per kilogram of direct materials is $10. The standard rate for labour is $16. Overhead is applied on the basis of direct labour hours. The variable overhead rate for maternity is $30 per hour, and the fixed overhead rate is $40 per hour.

Actual operating data for 2011 are as follows:

  1. Patient days produced: normal, 3,500; caesarean, 7,000.
  1. Direct materials purchased and used: 172,000 kilograms at $9.50-30,000 for normal maternity patients and 142,000 for the caesarean patients; no beginning or ending raw materials inventories.
  1. Nursing labour: 36,500 hours - 7,200 hours for normal patients and 29,300 hours for the caesarean; total cost of labour, $580,350.

Required:

  1. Prepare a standard cost sheet showing the unit cost per patient day for each type of patient.
  1. Compute the materials price and usage variances for each type of patient.
  1. Compute the labour rate and efficiency variances.
  1. Assume that you know only the total direct materials used for both procedures and the total direct labour hours used for both procedures. Can you compute the total materials usage and labour efficiency variances? Explain.
Answer #1

Answer :-

1. Normal patient day:

Standard Price Standard Usage

Standard Cost ( Standard Price × Standard Usage)

Direct Material $ 10 8 kg $ 80
Direct Labor $ 16 2 hrs $ 32
Variable Overhead $30 2 hrs $ 60
Fixed Overhead $ 40 2 hrs $ 80
Unit Cost (Total of Standard Cost) $ 252



Cesarean patient day:

Standard Price Standard Usage

Standard Cost ( Standard Price × Standard Usage)

Direct Material $ 10 20 kg $ 200
Direct Labor $ 16 4 hrs $ 64
Variable Overhead $30 4 hrs $ 120
Fixed Overhead $ 40 4 hrs

$ 160

Unit Cost (Total of Standard Cost) $ 544

Note :- Given ,Overhead is applied on the basis of direct labour hours.So Standard usage of Overhead is same as given Labor hours.


2. MPV= (AP – SP)AQ

Where, MPV = Material Price Variances

AP = Actual Price = $ 9.50

SP = Standard Price = $ 10

AQ = Actual Quantity

In Normal , AQ = 30,000

In Cesarean ,AQ = 142,000

In Normal,

MPV= ($9.50 – $10.00)30,000

MPV = $15,000 Favourable

In Cesarean

MPV = ($9.50 – $10.00)142,000

MPV = $71,000 Favourable

Now we find out Material Usage Variances
MUV = (AQ – SQ)SP

MUV = Material Usage Variances

AQ = Actual Quantity

In Normal , AQ = 30,000

In Cesarean ,AQ = 142,000

SQ = Standard Quantity = Direct Material Usage × Patient day produced

In Normal , SQ = 8 × 3,500 = 28,000

In Cesarean ,SQ = 20 × 7,000 = 140,000

In Normal

MUV = (30,000 – 28,000)$10

MUV = $20,000 Unfavorable

In Cesarean
MUV = (142,000 – 140,000 )$10

MUV = $20,000 Unfavorable


3. LRV = (AR – SR)AH

LRV = Labor rate Variances

AR = Actual rate = Total cost of Labor / Total Labor hours

AR = Actual rate = $580,350 / 36,500

AR = $15.90

SR = Standard Rate = $16

In Normal,
LRV = ($15.90 – $16.00)7,200

LRV = $720 Favourable

LRV= ($15.90 – $16.00)29,300

LRV = $2,930 Favourable

LEV = (AH – SH)SR

LRV = Labor Efficiency Variances

AH = Actual Hour

In Normal , AH = 7,200

In Cesarean ,AH = 29,300

SH = Standard Hours = Direct Labor hour× Patient day produced

In Normal , SH = 2× 3,500 = 7,000

In Cesarean ,AQ = 4× 7,000 = 28,000

SR = Standard Rate = $16

In Normal ,

LEV = (7,200 – 7000)$16

LEV = $3,200 Unfavorable

In Cesarean,
LEV = (29,300 – 28,000)$16

LEV = $20,800 Unfavorable


4. MUV= (AQ – SQ)SP

MUV = Material Usage Variances

AQ = Actual Quantity = 172,000

SQ = Standard Quantity of Normal and Cesarean = 28,000 + 140,000 = 168,000

SP = Standard Price = $10

MUV = (172,000 – 168,000)$10 = $40,000 Unfavorable

LEV = (AH – SH)SR

LRV = Labor Efficiency Variances

AH = Actual Hour = 36,500

Standard Rate = $16

LEV = (36,500 – 35,000)$16 = $24,000 U